OpenStudy (cheerleader):

Using the formula for finding the simple interest, I = Prt, find the interest earned on a savings account by depositing $9,800 for 15 months at 5% simple interest. $575.00 $612.50 $700.00 $642.50 None of these choices are correct.

OpenStudy (cheerleader):

@wolf1728

OpenStudy (wolf1728):

Interest = Principal * rate * time Do, you know the value of each?

OpenStudy (cheerleader):

9,800 for principal, 15 for time and 0.05 for interest

OpenStudy (wolf1728):

Very good!! and if you multiply all 3 9,800 * 15 * .05 you get?

OpenStudy (wolf1728):

Remember you are working with years.

OpenStudy (wolf1728):

So 15 should equal 15/12

OpenStudy (cheerleader):

22050 and 1.25

OpenStudy (wolf1728):

9,800 * 1.25 Years * .05 equals ?

OpenStudy (cheerleader):

612.5

OpenStudy (wolf1728):

That's right!!

OpenStudy (cheerleader):

Your really helpful

OpenStudy (wolf1728):

Okay so you have 2 more?

OpenStudy (cheerleader):

Yeaa

OpenStudy (cheerleader):

These are the answers: 10 % 11% 12.5% 9.5% None of these choices are correct.

OpenStudy (wolf1728):

You have to determine how $7,000 can become 14,000 in 8 years SO, to use that new formula, you would have to determine the rate for that to happen in 8 years. We already know the first "P" is 7,000 then the interest would be Int = 7,000 * r * 8 7,000 = 7,000 * 8 * r rate = 7,000 / (7,000 * 8) rate = 7,000 / 56,000 Can you solve that?

OpenStudy (cheerleader):

0.125

OpenStudy (wolf1728):

That's the rate and to show that as a percentage you would call it 12.5%

OpenStudy (cheerleader):

okay. so that is 12.5% * what

OpenStudy (wolf1728):

you want to know what per cent means?

OpenStudy (cheerleader):

yea

OpenStudy (wolf1728):

Per cent is Latin meaning "for every hundred". So, 12.5% rate means for every $100 you get $12.50

OpenStudy (cheerleader):

is it 612.50

OpenStudy (wolf1728):

No - you just solved it - the answer is 12.5%

OpenStudy (cheerleader):

oh okay thanks can i ask you one more?

OpenStudy (wolf1728):

I think you are looking at the answers for the SECOND question you asked

OpenStudy (wolf1728):

yes, one more

OpenStudy (cheerleader):

Yea i was

OpenStudy (wolf1728):

Okay

OpenStudy (cheerleader):

Algebraically manipulating the formula for finding the simple interest, find the principal (PV ) deposited in a savings account which earned $3,500 in interest at a simple interest rate of 8.5% for 10 years (to the nearest whole cent). $3,987.55 $4,117.65 $4,276.66 $4,550.48 None of these choices are correct.

OpenStudy (wolf1728):

Interest = Principal * rate * time They want us to solve for principal so the formula becomes Principal = Interest /(rate * time) Principal = 3,500 / (.085 * 10)

OpenStudy (cheerleader):

is it 4117.65

OpenStudy (wolf1728):

Yes that is it !!!

OpenStudy (cheerleader):

Omg thanks so much

OpenStudy (wolf1728):

u r welcome :-)